How many photons per second are emitted by a 50-kW radio transmitter that operates at a frequency of 1200 kHz?


1
Expert's answer
2021-12-21T12:16:42-0500

In one second, the transmitter uses the energy of


"E=Pt."

One electron has the following energy:


"E_e=hf."

The number of electrons:


"N=\frac E{E_e}=\frac{Pt}{hf},\\\space\\
N=\frac{50000·1}{(6.63·10^{-34})·(1200000)}=6.29·10^{31}."